# Mixture of Directed Graphical Models for Discrete Spatial Random Fields

### v3

@@ 6 Application to Survey Reports of Physical Disorder
In this section, we compare the MDGM-ST and the aMRF on an analysis of survey data from the Adolescent Health and Development in Context (AHDC) Study. This longitudinal study was designed to understand how social processes affect developmental outcomes of youth in the Columbus, Ohio metropolitan area. A representative sample of households residing within the I-270 belt loop and with youth ages 11-17 was obtained and a single youth from each household was randomly selected to participate in the study. Upon enrollment into the study, a caregiver of the youth, usually the mother, answered a questionnaire with a variety of questions about the conditions of their own neighborhood, as well as about locations which they routinely visit. We use data from the first wave of the study, which was collected between 2014-2016. One subset of questions – consistent with the ecometrics approach in urban sociology (Raudenbush and Sampson, 1999) – ask participants to report on aspects of the physical disorder of each of their routine activity locations, i.e. presence of garbage, graffiti, needles, liquor bottles, etc. For this analysis, we analyze the respondents’ perceptions of garbage, which was asked for neighborhoods as “In your neighborhood are [garbage, litter, broken glass] a big problem, somewhat of a problem or not a problem?” and for routine locations “Let us know if you consider [garbage, litter, broken glass] a big problem, somewhat of a problem, or not a problem at [location X]”. We recoded the three-scale Likert-type responses to a binary outcome with one corresponding to “a big problem” and “somewhat of a problem” and zero corresponding to “not a problem”. The spatial units of analysis are the 615 census block groups within the I-270 belt loop. The caregiver ratings are assumed to be conditionally independent within each block group. In total, there are 9469 ratings across the 615 block groups, of which 34 block groups have no ratings. Of the block groups that do have ratings, the median number of ratings is 9, the max is 192, and the third quartile is 19.
Figure 2: Left: map of the posterior mean of the latent process for each block group within the I-270 belt loop around Columbus, Ohio with a MDGM-ST prior. Center: The difference between the block group posterior mean of the latent variable for the MDGM-ST minus the aMRF prior. The block groups outlined in a thicker black line indicate units with no ratings. Right: Comparison of the posterior distributions for $\beta$ and $\boldsymbol{\eta}$ for the two different priors on the latent variable.
@@
We obtained 5000 draws from the posterior using the MCMC algorithms described in Section 4 and discarded the first 1000 as burnin. Figure 2 shows the posterior mean of the latent variable at each block group for an MDGM-ST prior in the left panel and the difference between the MDGM-ST minus the aMRF prior in the center panel. The difference between models is small and both produce a similar partition of the block groups into higher and lower probability areas of ratings indicating that garbage is a problem.
The center panel of Figure 2 shows the difference between the block group posterior means for the two models, with the block groups with zero ratings outlined in black. In the right panel of Figure 2 are box plots of the posterior distributions for the spatial dependence parameter $\beta$ and the Bernoulli error parameters $\boldsymbol{\eta}$. Unsurprisingly, the posterior distributions for $\beta$ are quite different in the two models. As the MDGM-ST class involves fewer connections between areal units in each update of the latent variable, a larger value of $\beta$ is needed to produce a similar degree of clustering of the latent variable as compared to an MRF. Overall, the MDGM-ST estimates slightly higher probabilities of the latent variable being one compared to the aMRF, but has lower estimates of the of error parameter. In Appendix E, we detail our comparison the performance of the two models through a cross-validation study by holding out data in training the model and then evaluating how well each model can predict the held out areal unit ratings. In the cross validation study, the performance between the MDGM-ST and aMRF priors is practically equivalent. The advantage of the MDGM-ST is the validity of the posterior distribution.
@@ 7 Related Work and Discussion
While we do not propose the MDGM framework as a formal approximation to an MRF model, we note that special cases of the framework appear in the literature as MRF approximation techniques. In particular, Cressie and Davidson (1998) proposed partially-ordered Markov models (POMMs) as a DGM-based alternative to an MRF model. Cressie and Davidson show that any POMM can be expressed equivalently as an MRF, and that for some MRFs, there is a POMM which closely approximates the probabilities of the configurations of a graph. We note that the DAG associated with the POMM is in the super set of DAGs compatible (as defined in Section 2.2) with the NUG associated with the MRF. In this way, POMMs are a special case of the MGDM framework in which there is only one component in the mixture. Other modeling strategies which use the computational advantages of DGMs include the the Markov mesh model (Abend et al., 1965), the POMM-inspired DGM approximation of the Potts model (the categorical extension of the two state Ising model) (Chakraborty et al., 2022), and a POMM-inspired approximation of the normalizing constant of an MRF (Tjelmeland and Austad, 2012).
@@
We also note the flexibility of the MDGM modeling framework. In particular, the parent conditional distributions, which determine the form of the DGM components in Equation 3, can be any valid probability distribution (e.g., Gaussian, Poisson). Additionally, as noted in the introduction, the MDGM can be used directly as a model for discrete outcomes, analogously to autologistic models where an MRF is used to directly model the data (Besag, 1974, 1975, 1986; Augustin et al., 1996; Hoeting et al., 2000; Hughes et al., 2011; Hughes, 2014; Caragea and Kaiser, 2009), instead of being used as a prior distribution in a hierarchical model as described in Section 3.
@@
Arguably, the standard class of spatial statistical models for categorical data, at least in the Bayesian setting, is one that uses the auxiliary variable formulation for probit regression of Albert and Chib (1993, 1997). The data augmentation strategy allows for the specification of a Gaussian MRF for a continuous latent variable, which is then related back to the discrete observed outcome by applying cuttoffs or “clipping” (De Oliveira, 1997; Higgs and Hoeting, 2010; Berrett and Calder, 2012, 2016; Schliep and Hoeting, 2015). When the Gaussian MRF is specificed using the conditional autoregressive (CAR) form (Besag, 1974), the model suffers from partial identifiability of the spatial dependence parameter (Carter et al., 2024). Alternatively, an intrinsic conditional autoregressive specification, which does not include a spatial dependence parameter for the covariance between areal units and, thus, avoids the identifiability problem. This specification corresponds, however, to an improper prior. While the posterior distribution of the latent spatial process is proper, the posterior predictive distribution for areal units without observed data is improper. As a consequence, the latent spatial process cannot be predicted at unobserved locations and cross-validation model assessments are limited. The MDGM framework, on the other hand, offers a proper posterior predictive distributions for settings with missing data and cross-validation-style model comparisons can readily be performed.
@@
In the continuously-indexed spatial data setting, we note that our work also parallels recent developments in scalable inference for Gaussian process (GP) models, for which there has been interest in exploiting the computational advantages of DGMs to fit GP models to massive datasets (Vecchia, 1988; Stein et al., 2004; Datta et al., 2016; Guinness, 2018; Katzfuss and Guinness, 2021; Katzfuss et al., 2020). Recent advances include the meshed GPs of Peruzzi et al. (2022), which assigns a DAG to the subsets of a partition of the domain; a DGM approximation of Nearest Neighbor GPs to improve computation of spatial probit linear mixed models (Saha et al., 2022); and a “Bag of DAGs” model to improve computational speed and learn prevailing wind patterns in the spread of air pollutants (Jin et al., 2023).
In conclusion, the MDGM is a flexible and computationally efficient framework for modeling observations associated with a spatial areal units. Most importantly for the discrete outcome setting, the MDGM offers a computationally fast and simple implementation alternative to the widely used, yet theoretically tenuous, aMRF. Additionally in the variety of data settings described above, the MDGM offers a flexible alternative to standard models.
Beyond applying the MDGM to different data settings (e.g., continuous outcomes, multicategory outcomes), in future work we will explore alternative MCMC schemes – as an alternative to our independent MH proposals – for the AO and rooted MDGM classes. Through our simulation studies, we found the ST class of MGDMs performed particularly well in estimating the latent field, both in terms of accuracy and capturing the spatial dependence of the field. The minimally connected ST-DAGs appear to offer greater flexibility in modeling the spatial dependence across areal units compared to the AO and rooted classes when the data are generated from an underlying MRF. We will explore whether this result holds for other data generating mechanisms.
@@
Includes the proof of Proposition 3.1, discussion of concepts and examples related to the MDGM, and additional simulation study results. (.pdf file)
@@
The anonymized ratings of levels of garbage for the block groups within the I-270 belt loop of Columbus, Ohio from the AHDC Study. The file also includes the spatial features data frame for the block groups and corresponding second-order NUG. (.Rbin file)
@@ A.2 Examples of Compatible and Incompatible DAGs
Figure 4 shows examples of DAGs that are compatible and incompatible with the first-order NUG shown in the top left of the figure. The top center DAG is an example of a directed spanning tree which is compatible with the NUG. The acyclic orientation DAG in the top right is also compatible with the NUG. The bottom row shows DAGs that are incompatible with the NUG as they introduce directed edges between vertices which do not appear as undirected edges in the NUG.

### v4

@@ 6 Application to Survey Reports of Physical Disorder
In this section, we compare the MDGM-ST and the aMRF on an analysis of survey data from the Adolescent Health and Development in Context (AHDC) Study. This longitudinal study was designed to understand how social processes affect developmental outcomes of youth in the Columbus, Ohio metropolitan area. A representative sample of households residing within the I-270 belt loop and with youth ages 11-17 was obtained and a single youth from each household was randomly selected to participate in the study. Upon enrollment into the study, a caregiver of the youth, usually the mother, answered a questionnaire with a variety of questions about the conditions of their own neighborhood, as well as about locations which they routinely visit. We use data from the first wave of the study, which was collected between 2014-2016. One subset of questions – consistent with the ecometrics approach in urban sociology (Raudenbush and Sampson, 1999) – ask participants to report on aspects of the physical disorder of each of their routine activity locations, i.e. presence of garbage, graffiti, needles, liquor bottles, etc. For this analysis, we analyze the respondents’ perceptions of garbage, which was asked for neighborhoods as “In your neighborhood are [garbage, litter, broken glass] a big problem, somewhat of a problem or not a problem?” and for routine locations “Let us know if you consider [garbage, litter, broken glass] a big problem, somewhat of a problem, or not a problem at [location X]”. We recoded the three-scale Likert-type responses to a binary outcome with one corresponding to “a big problem” and “somewhat of a problem” and zero corresponding to “not a problem”. The spatial units of analysis are the 615 census block groups within the I-270 belt loop. We specify a first-order dependence structure for the NUG in this analyis. The caregiver ratings are assumed to be conditionally independent within each block group. In total, there are 9469 ratings across the 615 block groups, of which 34 block groups have no ratings. Of the block groups that do have ratings, the median number of ratings is 9, the max is 192, and the third quartile is 19.
Figure 2: Left: map of the posterior mean of the latent process for each block group within the I-270 belt loop around Columbus, Ohio with a MDGM-ST prior. Center: The difference between the block group posterior mean of the latent variable for the MDGM-ST minus the aMRF prior. The block groups outlined in a thicker black line indicate units with no ratings. Right: Comparison of the posterior distributions for $\boldsymbol{\eta}$ and $T(\mathbf{z})/M$, where $M=|\bar{\mathbf{E}}|$ for the two different priors on the latent variable.
@@
We ran four MCMC chains in parallel for each model, with 5000 draws per chain, using the MCMC algorithms described in Section 4 and Appendix E. The latent variable was independently and random initiated. We initiated $\eta_{0}$ with random uniform draw on $(0,.5)$, $\eta_{1}$ on $(0.5,1)$, and $\beta$ on $(0,0.88)$ for the aMRF and $(0,2)$ for the MDGM-ST. We discarded the first 1000 draws of each chain as burnin. The estimated R-hat values for each parameter and the statistic $T(\mathbf{z})$ were all less than 1.01 and trace plots of the four chains for each parameter showed good mixing.
Figure 2 shows the posterior mean of the latent variable at each block group for an MDGM-ST prior in the left panel and the difference between the MDGM-ST minus the aMRF prior in the center panel. The difference between models is small and both produce a similar partition of the block groups into higher and lower probability areas of ratings indicating that garbage is a problem. Block groups with zero ratings outlined in black.
In the right panel of Figure 2 are box plots of the posterior distributions for the Bernoulli error parameters $\boldsymbol{\eta}$ and $T(\mathbf{z})/|\bar{\mathbf{E}}|$. Overall, the MDGM-ST estimates slightly higher probabilities of the latent variable being one compared to the aMRF, but has lower estimates of the error parameter. Additionally, the MDGM-ST estimates a higher proportion of matching neighboring block groups with respect to the NUG than the aMRF. In Appendix I, we detail our comparison the performance of the two models through a cross-validation study by holding out data in training the model and then evaluating how well each model can predict the held out areal unit ratings. In the cross validation study, the performance between the MDGM-ST and aMRF priors is practically equivalent. The advantage of the MDGM-ST is the validity of the posterior distribution.
@@ 7 Related Work and Discussion
While we do not propose the MDGM framework as a formal approximation to an MRF model, we note that special cases of the framework appear in the literature as MRF approximation techniques. In particular, Cressie and Davidson (1998) proposed partially-ordered Markov models (POMMs) as a DGM-based alternative to an MRF model. Cressie and Davidson show that any POMM can be expressed equivalently as an MRF, and that for some MRFs, there is a POMM which closely approximates the probabilities of the configurations of a graph. We note that the DAG associated with the POMM is in the set of DAGs compatible (as defined in Section 2) with the NUG associated with the MRF. In this way, POMMs are a special case of the MGDM framework in which there is only one component in the mixture. Other modeling strategies which use the computational advantages of DGMs include the Markov mesh model (Abend et al., 1965), the POMM-inspired DGM approximation of the Potts model (the categorical extension of the two state Ising model) (Chakraborty et al., 2022), and a POMM-inspired approximation of the normalizing constant of an MRF (Tjelmeland and Austad, 2012).
@@
We also note the flexibility of the MDGM modeling framework. In particular, the parent conditional distributions, which determine the form of the DGM components in Equation 3, can be any valid probability distribution (e.g., Gaussian, Poisson). Additionally, as noted in the introduction, the MDGM can be used directly as a model for discrete outcomes, analogously to autologistic models where an MRF is used to directly model the data (Besag, 1974, 1975, 1986; Augustin et al., 1996; Hoeting et al., 2000; Hughes et al., 2011; Hughes, 2014; Caragea and Kaiser, 2009), instead of being used as a prior distribution in a hierarchical model as described in Section 3. Lastly, the spanning tree class provides a more flexible characterization of spatial dependence through a global spatial dependence parameter and local adjustments from posterior edge weights.
@@
Arguably, the standard class of spatial statistical models for categorical data, at least in the Bayesian setting, is one that uses the auxiliary variable formulation for probit regression of Albert and Chib (1993, 1997). The data augmentation strategy allows for the specification of a Gaussian MRF for a continuous latent variable, which is then related back to the discrete observed outcome by applying cuttoffs or “clipping” (De Oliveira, 1997; Higgs and Hoeting, 2010; Berrett and Calder, 2012, 2016; Schliep and Hoeting, 2015). When the Gaussian MRF is specificed using the conditional autoregressive (CAR) form (Besag, 1974), the model suffers from partial identifiability of the spatial dependence parameter (Carter et al., 2024). Alternatively, an intrinsic conditional autoregressive specification, which does not include a spatial dependence parameter for the covariance between areal units and, thus, avoids the identifiability problem. This specification corresponds, however, to an improper prior. While the posterior distribution of the latent spatial process is proper, the posterior predictive distribution for areal units without observed data is improper. As a consequence, the latent spatial process cannot be predicted at unobserved locations and cross-validation model assessments are limited. The MDGM framework, on the other hand, offers proper posterior predictive distributions for settings with missing data and cross-validation-style model comparisons can readily be performed.
@@
In the continuously-indexed spatial data setting, we note that our work also parallels recent developments in scalable inference for Gaussian process (GP) models, for which there has been interest in exploiting the computational advantages of DGMs to fit GP models to massive datasets (Vecchia, 1988; Stein et al., 2004; Datta et al., 2016; Guinness, 2018; Katzfuss and Guinness, 2021; Katzfuss et al., 2020; Dey et al., 2022). Recent advances include the meshed GP of Peruzzi et al. (2022), which assigns a DAG to the subsets of a partition of the domain; a DGM approximation of Nearest Neighbor GPs to improve computation of spatial probit linear mixed models (Saha et al., 2022); and a “Bag of DAGs” model to improve computational speed and learn prevailing wind patterns in the spread of air pollutants (Jin et al., 2023).
In conclusion, the MDGM is a flexible and computationally efficient framework for modeling observations associated with spatial areal units. Most importantly for the discrete outcome setting, the MDGM offers a computationally fast and simple implementation alternative to the widely used, yet theoretically tenuous, aMRF. Additionally, in the variety of data settings described above, the MDGM offers a flexible alternative to standard models.
Beyond applying the MDGM to different data settings (e.g., continuous outcomes, multicategory outcomes), in future work we will explore alternative MCMC schemes – as an alternative to our independent MH proposals – for the acyclic orientation and rooted MDGM classes. Through our simulation studies, we found the spanning tree class of MGDMs performed particularly well in estimating the latent field, both in terms of accuracy and capturing the spatial dependence of the field. In terms of approximation, the strong compatibility of spanning trees provides a probable theoretical reason why the class comes closest to the posterior with an MRF prior in our simulation studies. The minimally connected ST-DAGs appear to offer greater flexibility in modeling the spatial dependence across areal units compared to the acyclic orientation class when the data are generated from an underlying MRF. We will explore whether this result holds for other data generating mechanisms.
@@
Includes all proofs, discussion of concepts and examples related to the MDGM, two additional MDGM classes, and additional simulation study results. (.pdf file)
@@
The anonymized ratings of levels of garbage for the block groups within the I-270 belt loop of Columbus, Ohio from the AHDC Study. The file also includes the spatial features data frame for the block groups and corresponding first-order NUG. (.Rbin file)
@@ A.2 Examples of Compatible and Incompatible DAGs
Figure 4 shows examples of DAGs that are compatible and incompatible with the first-order NUG shown in the top left of the figure. The top center DAG is an example of a directed spanning tree which is strongly compatible with the NUG. The acyclic orientation DAG in the top right is weakly compatible with the NUG. The bottom row shows DAGs that are incompatible with the NUG as they introduce directed edges between vertices which do not appear as undirected edges in the NUG.